# CoreMark: Toward Robust and Universal Text Watermarking Technique

### v1

@@ IV-A Watermark Embedding
Let $C=\left\{c_{i},i=1,...,n_{c}\right\}$ be the set of all characters in the text image, and $L_{\text{core}}=\left\{l_{\text{core}}^{i},i=1,...,n_{c}\right\}$ be the set of lengths of all $S_{\text{core}}$s in $C$, where $n_{c}$ is the number of all characters. Each $l_{i}$ can be calculated using Eq. 6:
$l_{\text{core}}^{i}=\frac{\sum_{j=1}^{\text{len($\mathcal{I}$)}}\mu_{j}^{\text{max}}}{\text{len($\mathcal{I}$)}}.$ (6)
We establish a threshold $T_{\lambda}$ to determine character eligibility. Notably, $T_{\lambda}$ is not fixed but calculated based on the lengths of $S_{\text{core}}$s in the document. To calculate $T_{\lambda}$, we first sort all lengths of $S_{\text{core}}$s from longest to shortest according to $L_{\text{core}}$, with the sorted values represented by $\tilde{L}_{\text{core}}=\left\{\tilde{l}_{\text{core}}^{i},i=1,...,n_{c}\right\}$. Assuming $\lambda$ is a value less than 1, we calculate $T_{\lambda}$ as follows:
$T_{\lambda}=\tilde{l}_{\text{core}}^{\mathop{\min}\left\{i\mid\tilde{l}_{i}>\tilde{l}_{i\times\lambda}\right\}}.$ (7)
Let W = $w_{i},i=1,...,L_{w}$ be a watermark sequence with length $L_{w}$ to be embedded. We first assess whether the length of $S_{\text{core}}$ for the current character $c_{i}$ exceeds $T_{\lambda}$; If it does, we embed one bit of information in $c_{i}$; if not, we apply the same evaluation to the next character. This process continues until all information has been embedded.
@@
Component 3: Embedding Mechanism. In our proposed scheme, the thickness of $S_{\text{core}}$, defined by the number of ‘candidate vectors’ it contains, is used as the embedding feature for watermarking. To embed a bit “0”, we ensure the thickness of $S_{\text{core}}$ exceeds a predetermined threshold $T_{\Delta}$ by $t$. Conversely, to embed a bit ”1”, we constrain the thickness of $S_{\text{core}}$ to be less than $T_{\Delta}$ by $t$. Here, $t$ represents the watermark embedding strength. Let $\alpha_{i}$ and $\tilde{\alpha}_{i}$ denote the thickness of $S_{\text{core}}$ before and after embedding, respectively, and let $w_{i}$ represent the watermark to be embedded. The watermark embedding process for each character can then be expressed mathematically as:
${\tilde{\alpha_{i}}}=\begin{cases}T_{\Delta}-t,&\text{if $\Delta\leq t$ and $w_{i}=1$}\\ \alpha_{i},&\text{if $\Delta>t$ and $w_{i}=1$}\\ T_{\Delta}+t,&\text{if $\Delta\geq t$ and $w_{i}=0$}\\ \alpha_{i},&\text{if $\Delta<t$ and $w_{i}=0$}\\ \end{cases}.$ (8)
Specifically, we select a specific text line as a benchmark for calculating $T_{\Delta}$. Characters in this benchmark line are excluded from watermark embedding. If $N_{q}$ represents the number of characters in the benchmark line, $T_{\Delta}$ can be formulated as:
$T_{\Delta}=\frac{1}{N_{q}}\sum_{i=1}^{N_{q}}{\alpha_{i}}.$ (9)
In addition, $\Delta$ is computed as $\Delta=T_{\Delta}-\alpha_{i}$.
We satisfy the condition specified in Eq. 8 by either thinning or thickening $S_{\text{core}}$. In particular, human perceptual factors are considered by analyzing neighboring pixels of $S_{\text{core}}$ to ensure that modifications do not introduce visually perceptible differences. The methods for thickening and thinning are discussed separately below.
IV-A1 Thinning
To ensure that $C_{\text{max}}^{\text{ave}}$ remains the largest value among $C_{1,...,k}^{\text{ave}}$ after thinning, we select the shorter ‘candidate vectors’ on both sides of $S_{\text{core}}$ and convert their black pixels to white. Using the horizontal direction as an example, the thinning process follows these steps:
@@ 1. Step 1:
    Compare $l_{1}$ and $l_{n}$ to identify the shorter side.
@@ 2. Step 2:
    If $l_{1}$ is shorter, use row $\mathcal{I}_{1}\!-\!1$ as a reference. A black pixel in row $\mathcal{I}_{1}$ is deemed ‘flippable’ if the adjacent pixel in row $\mathcal{I}_{1}\!-\!1$ is white. Conversely, if $l_{n}$ is shorter, use row $\mathcal{I}_{\text{len($\mathcal{I}$)}}+1$ as the reference row.
@@ 3. Step 3:
    Repeat the operation in Step 2 for $\text{abs}(\Delta)+1$ iterations to complete the thinning process.
IV-A2 Thickening
Contrast to thinning, we flip the pixels that do not belong to $S_{\text{core}}$ to thicken a $S_{\text{core}}$ so that the thickness of $S_{\text{core}}$ is larger than $T_{\Delta}$ after thickening. The process of thickening is listed as follows:
@@ 1. Step 1:
    Determine the longer side by comparing $l_{1}$ and $l_{n}$.
@@ 2. Step 2:
    If $l_{1}$ is greater, flip all white pixels to black within the region bounded by abscissas from $\mu_{\mathcal{I}_{1}}^{\text{max}}-\text{abs}(\Delta)-1$ to $\mu_{\mathcal{I}_{1}}^{\text{max}}-1$ and ordinating from $y_{\mu_{\mathcal{I}_{1}}^{\text{max}}}$ to $y_{\mu_{\mathcal{I}_{1}}^{\text{max}}}+\mu_{\mathcal{I}_{1}}^{\text{max}}$. Alternatively, if $l_{n}$ is greater, flip all white pixels to black within the region bounded by abscissas from $\mu_{\mathcal{I}_{\text{len($\mathcal{I}$)}}}^{\text{max}}+1$ to $\mu_{\mathcal{I}_{\text{len($\mathcal{I}$)}}}^{\text{max}}+\text{abs}(\Delta)+1$ and ordinating from $y_{\mu_{\mathcal{I}_{\text{len($\mathcal{I}$)}}}^{\text{max}}}$ to $y_{\mu_{\mathcal{I}_{\text{len($\mathcal{I}$)}}}^{\text{max}}}+\mu_{\mathcal{I}_{\text{len($\mathcal{I}$)}}}^{\text{max}}$.
@@ IV-B Watermark Extraction
To extract the hidden message from a text image, we first segment all characters and extract $S_{\text{core}}$s of each one. Next, robust characters are selected as described in Section 7. Let $\alpha_{i}^{{}^{\prime}}$ represent the thickness of $S_{\text{core}}$ in the extraction stage. After computing $T_{\Delta}^{{}^{\prime}}$ by Eq 9, we can extract the watermark by comparing $\alpha_{i}^{{}^{\prime}}$ with $T_{\Delta}^{{}^{\prime}}$. The watermark detection can be formulated as follows.
$w_{i}=\begin{cases}1,&\text{if $T_{\Delta}^{{}^{\prime}}-\alpha_{i}\geq 0$}\\ 0,&\text{if $T_{\Delta}^{{}^{\prime}}-\alpha_{i}<0$}\\ \end{cases}.$ (10)
Figure 9: Data hiding in English text document image and Chinese text document image: (a) original copy (Chinese); (b) marked copy (Chinese); (c) original copy (English); (d) marked copy (English); (e1, f1) magnified partial of (a); (e2) difference between (e1) and (e3) (shown in black); (e3) magnified partial of (b) (after thinning); (f2) difference between (f1) and (f3) (shown in black); (f3) magnified partial of (b) (after thickening); (g1, h1) magnified partial of (c); (g2) difference between (g1) and (g3) (shown in black); (g3) magnified partial of (c) (after thinning); (h2) difference between (h1) and (h3) (shown in black); (h3) magnified partial of (d) (after thickening);
@@ V-B Main Results
Watermark Imperceptibility. Figure 9(a) and Figure 9(c) show the original Chinese and English documents, respectively. Figure 9(b) and Figure 9(d) are the corresponding watermarked documents. On the right side of Figure 9, we enlarge portions of the original images and the marked images and represent their differences with black pixels. We can see that some characters are not modified, such as the second character “,” in Figure 9(e2) and the first character “n” in Figure 9(h2). The reasons are: 1) The character was not selected for embedding, as explained in Section 7; 2) $\Delta>0$ when embedding “1”, and $\Delta<0$ when embedding “0”, as described in Eq. 8, respectively. Table I reports a quantitative comparison of visual quality between original and watermarked images across a variety of font styles. The proposed CoreMark consistently outperforms three baselines across both English and Chinese font families. Specifically, CoreMark achieves the highest PSNR and SSIM scores in all cases, with PSNR gains of up to 11.31 dB and SSIM values exceeding 0.999 in most cases, indicating minimal visual distortion. These results highlight the superior imperceptibility of CoreMark across a wide range of font styles and demonstrate its effectiveness in preserving visual quality.
TABLE I: Robustness ($ACC(\%)$) against screenshots of text with different sizes in different languages.
@@
Watermark Robustness (against Screenshots). Text images are often propagated via screenshots, with distortions primarily arising from the down-sampling of the screen. As shown in Table II, no matter Chinese fonts or English fonts, all $ACC$s of our CoreMark exceed 94%, which significantly outperform the baseline methods. Even in the case of 10pt, the extraction success rates of CoreMark for Chinese and English fonts are still above 94% and 95%, respectively, which indicates that CoreMark can achieve satisfactory robustness even with smaller fontsizes. In contrast, the performance of baseline methods exhibits noticeable limitations across both Chinese and English texts. Wu04 and AutoStegaFont yield relatively low extraction accuracy, with their lack of robustness analyzed in detail in Sections III-B and III-D, respectively. Although Yang23 demonstrates improved resilience against screenshots, achieving accuracies ranging from 78.61% to 95.43% for Chinese and 81.37% to 93.83% for English, its performance remains unstable, particularly under smaller font sizes.
TABLE III: Comparison results of the robustness of CoreMark without and with IES modulator in the printing and scanning process under different device combinations.
@@
TABLE IV: Qualitative Analysis of CoreMark in Different Fonts.
@@
TABLE V: Qualitative Analysis of CoreMark in Foreign Languages.
@@
Watermark Robustness (against Print-scanning). Printers and scanners are widely used in daily document processing due to their accessibility and convenience. Printing generates physical document copies while scanning digitizes hardcopy materials for electronic storage, editing, or analysis. To verify the robustness against print-scan, we first print text images with different fontsizes on A4 paper using HP LaserJet Pro MFP M132nw at 600 dpi, then digitize them using Epson V30 SE scanner at 600 dpi. As shown in Table IV, our approach consistently achieves $>93$% accuracy at 10pt for both Chinese and English texts, significantly outperforming baseline methods, particularly at smaller font sizes. To assess cross-device generalizability, we test CoreMark across multiple printer-scanner combinations using 50 randomly selected images with varied font sizes. Table III presents the extraction accuracy across two printer models at 600dpi (HP LaserJet Pro MFP M132nw and HP LaserJet Pro M17w) and three scanners at 600 dpi (Epson V30, M132nw, and Canon LiDE400). CoreMark shows strong robustness across all configurations, with further improvements when integrated with the ES modulator, confirming the effectiveness of the ES modulator in enhancing resilience against cross-device variations.
TABLE VI: Average Extraction Accuracy Under Varying Shooting Distances and Angles.
@@
96.28 97.69 97.24 95.27 94.61 97.67 95.33 95.00 94.674 93.67
Watermark Robustness (against Print-camera). Print-camera shooting is also a common method for capturing text content. We first print all the encoded text images on A4 paper at 600 dpi using the printer HP M132nw and take photos with a handheld cellphone HONOR 30S. The shooting distance is 20cm and the shooting angle is $0^{\circ}$. Subsequently, watermark extraction is performed on the images. As shown in Table IV, CoreMark maintains stable high accuracy, clearly outperforming the best competing baseline, Yang23 [9]. When the Chinese font size exceeds 16pt, the $ACC$ of Yang23 surpasses 90%. But when the font size is smaller, its extraction accuracy drops sharply. To further assess the robustness of CoreMark against print-camera, we also evaluate it under other shooting conditions. The results, shown in Table VI, demonstrate consistently high accuracy across all conditions, confirming the strong robustness of CoreMark against common real-world perturbations such as distance variation and angular deviation during image capture.
@@
Watermark Generalizability (w.r.t. Different Fonts). We evaluate the robustness and imperceptibility of several commonly used fonts in both Chinese and English. The font sizes for Chinese and English fonts are set to 14pt and 18pt, respectively. For each font, we test 30 text images, and Table IV records the detailed results. Under the Screenshot, CoreMark attains near-perfect accuracy, reaching up to 99.39% on English (TNR) and 98.19% on Chinese (DengXian). Under more challenging conditions such as Print-scan and Print-camera, CoreMark maintains robust performance, demonstrating its superior resilience to physical distortions. Yang23 performs better than Wu04, particularly on English fonts, but still falls short compared to CoreMark. These results validate the generalizability of CoreMark across diverse font styles.
Watermark Generalizability (w.r.t. Foreign Languages). We also evaluate the adaptability of CoreMark to languages beyond English and Chinese, including Arabic, French, Japanese, Korean, Russian, and Spanish. For each language, we test 30 text images. Table V presents the PSNR, SSIM, and $ACC$ values of six languages. The results indicate that Arabic exhibits slightly lower robustness compared to other languages, but it achieves over 90% accuracy in print-scan and screenshot scenarios, and over 85% accuracy in resisting print-camera. Regardless of invisibility or robustness, CoreMark performs significantly better than the baseline methods.
Figure 13: Thinning examples in (a and b) Arabic, (c) and (d) Japanese, (e) and (f) Korean, (g) and (h) Russian.

### v2

@@ IV-A Watermark Embedding
Recall that $\mathcal{L}=\{\ell_{i}\}_{i=1}^{n_{\text{core}}}$ recordes the length of each sequence in $S_{\text{core}}$. The overall length of $S_{\text{core}}$, denoted as $\ell^{\text{core}}$, is the average length across all sequences and is defined below:
$\ell^{\text{core}}=\frac{\sum_{i=1}^{n_{\text{core}}}\ell_{i}}{n_{\text{core}}},$ (6)
Let $C=\{c_{i}\}_{i=1}^{n_{c}}$ be the set of all characters in a text image, where $n_{c}$ is the total number of characters contained in the image. The core lengths of all characters are stored in the set $L_{\text{core}}=\{\ell^{\text{core}}_{i}\}_{i=1}^{n_{c}}$. We establish a threshold $T_{\lambda}$ to determine character eligibility. Notably, $T_{\lambda}$ is not fixed but calculated based on the lengths of $S_{\text{core}}$s in the document. To calculate $T_{\lambda}$, we first sort all lengths of $S_{\text{core}}$s from shortest to longest according to $L_{\text{core}}$, with the sorted values represented by $\tilde{L}_{\text{core}}=\{\tilde{\ell}^{\text{core}}_{i}\}_{i=1}^{n_{c}}$. Assuming $\lambda\in(0,1]$ is a user-chosen percentile, we calculate $T_{\lambda}$ as follows:
$k_{\lambda}=\lceil n_{c}\cdot\lambda\rceil,\quad T_{\lambda}=\tilde{\ell}^{\text{core}}_{k_{\lambda}}.$ (7)
We set $\lambda=0.2$ in this paper.
Let W = $\{w_{i}\}_{i=1}^{L_{w}}$ be a watermark sequence with length $L_{w}$ to be embedded. We first assess whether the length of $S_{\text{core}}$ for the current character $c_{i}$ exceeds $T_{\lambda}$; If it does, we embed one bit of information in $c_{i}$; if not, we apply the same evaluation to the next character. This process continues until all information has been embedded.
@@
Component 3: Embedding Mechanism. In our proposed scheme, the embedding feature for watermarking is defined as the size of $S_{\text{core}}$, i.e., $n_{\text{core}}$. To embed a bit “0”, we ensure $n_{\text{core}}$ exceeds a predetermined threshold $T_{\Delta}$ by $\beta$. Conversely, to embed a bit ”1”, we constrain $n_{\text{core}}$ to be less than $T_{\Delta}$ by $\beta$. Here, $\beta$ represents the watermark embedding strength. Let $\tilde{n}_{\text{core}}$ denote the thickness of $S_{\text{core}}$ after embedding. The watermark embedding process for each character can then be expressed mathematically as:
[…]
Specifically, we designate a specific text line as the baseline for calculating the threshold $T_{\Delta}$, defined as the mean of $n_{\text{core}}$ values over all characters in this line. Characters within this baseline are not subject to watermark embedding. We satisfy the condition specified in Eq. 8 by either reducing or expanding $S_{\text{core}}$. In particular, human perceptual factors are considered by analyzing neighboring pixels of $S_{\text{core}}$ to ensure that modifications do not introduce visually perceptible differences. The methods for reducing or expanding are discussed separately below.
IV-A1 Reducing $S_{\text{core}}$
To ensure that $C_{\max}^{\text{ave}}$ remains the largest value among $C_{\text{ave}}$ after reducing, we select the shorter candidate vectors on both sides of $S_{\text{core}}$ and convert their black pixels to white. Using the horizontal direction as an example, the reducing process follows these steps:
@@ 1. Step 1:
    Compare $\ell_{1}$ and $\ell_{n_{\text{core}}}$ to identify the shorter side. Let $k$ denote the index of the shorter segment, i.e., $k=1$ if $\ell_{1}$ is shorter and $k=n_{\text{core}}$ if $\ell_{n_{\text{core}}}$ is shorter.
@@ 2. Step 2:
    Use row $\mathcal{I}_{k+d}$ as the reference, where d=1d=-1 if $k=1$ and $d=+1$ if $k=n_{\text{core}}$. A black pixel in row $\mathcal{I}_{k}$ is deemed ‘flippable’ if the adjacent pixel in the reference row $\mathcal{I}_{k+d}$ is white. After flipping, set $k\leftarrow k+1$ if $k=1$; set $k\leftarrow k-1$ if $k=n_{\text{core}}$.
@@ 3. Step 3:
    Repeat the operation in Step 2 for $|T_{\Delta}-n_{\text{core}}|+1$ iterations to complete the reducing process.
IV-A2 Expanding $S_{\text{core}}$
In contrast to reducing, pixels outside $S_{\text{core}}$ are converted to black in order to expand $S_{\text{core}}$. The process of expanding is listed as follows:
@@ 1. Step 1:
    Determine the longer side by comparing $\ell_{1}$ and $\ell_{n_{\text{core}}}$.
@@ 2. Step 2:
    If $\ell_{1}$ is longer, flip all white pixels to black within the region bounded by x-coordinates ranging from $\mathcal{I}_{1}-|T_{\Delta}-n_{\text{core}}|-1$ to $\mathcal{I}_{1}-1$ and y-coordinates ranging from $y_{\mathcal{I}_{1}}$ to $y_{\mathcal{I}_{1}}+\ell_{1}$. Alternatively, if $\ell_{n_{\text{core}}}$ is greater, flip all white pixels to black within the region bounded by x-coordinates ranging from $\mathcal{I}_{n_{\text{core}}}+1$ to $\mathcal{I}_{n_{\text{core}}}+|T_{\Delta}-n_{\text{core}}|+1$ and y-coordinates ranging from $y_{\mathcal{I}_{n_{\text{core}}}}$ to $y_{\mathcal{I}_{n_{\text{core}}}}+\ell_{\mathcal{I}_{n_{\text{core}}}}$.
@@ IV-B Watermark Extraction
To extract the hidden message from a text image, we first segment all characters and extract $S_{\text{core}}$s of each one. Next, robust characters are selected as described in Section 7. In the extraction stage, let $\hat{T}_{\Delta}$ denote the extraction threshold and $\hat{n}_{\text{core}}$ denote the size of $S_{\text{core}}$. The watermark bit can be extracted by comparing $\hat{n}_{\text{core}}$ with $\hat{T}_{\Delta}$. The watermark detection process can be formulated as follows:
$w_{i}=\begin{cases}1,&\text{if $\hat{T}_{\Delta}-\hat{n}_{\text{core}}\geq 0$}\\ 0,&\text{if $\hat{T}_{\Delta}-\hat{n}_{\text{core}}<0$}\\ \end{cases}.$ (9)
Figure 9: Data hiding in English text document image and Chinese text document image: (a) original copy (Chinese); (b) marked copy (Chinese); (c) original copy (English); (d) marked copy (English); (e1, f1) magnified partial of (a); (e2) difference between (e1) and (e3) (shown in black); (e3) magnified partial of (b) (after reducing $S_{\text{core}}$); (f2) difference between (f1) and (f3) (shown in black); (f3) magnified partial of (b) (after expanding $S_{\text{core}}$); (g1, h1) magnified partial of (c); (g2) difference between (g1) and (g3) (shown in black); (g3) magnified partial of (c) (after reducing $S_{\text{core}}$); (h2) difference between (h1) and (h3) (shown in black); (h3) magnified partial of (d) (after expanding $S_{\text{core}}$);
@@ V-B Main Results
Watermark Imperceptibility. Figure 9(a) and Figure 9(c) show the original Chinese and English documents, respectively. Figure 9(b) and Figure 9(d) are the corresponding watermarked documents. On the right side of Figure 9, we enlarge portions of the original images and the marked images and represent their differences with black pixels. We can see that some characters are not modified, such as the second character “,” in Figure 9(e2) and the first character “n” in Figure 9(h2). The reasons are: 1) The character was not selected for embedding, as explained in Section 7; 2) $\Delta>0$ when embedding “1”, and $\Delta<0$ when embedding “0”, as described in Eq. 8, respectively. Table II reports a quantitative comparison of visual quality between original and watermarked images across a variety of font styles. The proposed CoreMark consistently outperforms three baselines across both English and Chinese font families. Specifically, CoreMark achieves the highest PSNR and SSIM scores in all cases, with PSNR gains of up to 11.31 dB and SSIM values exceeding 0.999 in most cases, indicating minimal visual distortion. These results highlight the superior imperceptibility of CoreMark across a wide range of font styles and demonstrate its effectiveness in preserving visual quality.
@@
TABLE III: Robustness ($ACC(\%)$) against screenshots of text with different sizes in different languages.
@@
Watermark Robustness (against Screenshots). Text images are often propagated via screenshots, with distortions primarily arising from the down-sampling of the screen. As shown in Table III, no matter Chinese fonts or English fonts, all $ACC$s of our CoreMark exceed 94%, which significantly outperform the baseline methods. Even in the case of 10pt, the extraction success rates of CoreMark for Chinese and English fonts are still above 94% and 95%, respectively, which indicates that CoreMark can achieve satisfactory robustness even with smaller fontsizes. In contrast, the performance of baseline methods exhibits noticeable limitations across both Chinese and English texts. Wu04 and AutoStegaFont yield relatively low extraction accuracy, with their lack of robustness analyzed in detail in Sections III-B and III-D, respectively. Although Yang23 demonstrates improved resilience against screenshots, achieving accuracies ranging from 78.61% to 95.43% for Chinese and 81.37% to 93.83% for English, its performance remains unstable, particularly under smaller font sizes.
TABLE IV: Comparison results of the robustness of CoreMark without and with IES modulator in the printing and scanning process under different device combinations.
@@
TABLE V: Qualitative Analysis of CoreMark in Different Fonts.
@@
TABLE VI: Qualitative Analysis of CoreMark in Foreign Languages.
@@
Watermark Robustness (against Print-scanning). Printers and scanners are widely used in daily document processing due to their accessibility and convenience. Printing generates physical document copies while scanning digitizes hardcopy materials for electronic storage, editing, or analysis. To verify the robustness against print-scan, we first print text images with different fontsizes on A4 paper using HP LaserJet Pro MFP M132nw at 600 dpi, then digitize them using Epson V30 SE scanner at 600 dpi. As shown in Table V, our approach consistently achieves >93>93% accuracy at 10pt for both Chinese and English texts, significantly outperforming baseline methods, particularly at smaller font sizes. To assess cross-device generalizability, we test CoreMark across multiple printer-scanner combinations using 50 randomly selected images with varied font sizes. Table IV presents the extraction accuracy across two printer models at 600dpi (HP LaserJet Pro MFP M132nw and HP LaserJet Pro M17w) and three scanners at 600 dpi (Epson V30, M132nw, and Canon LiDE400). CoreMark shows strong robustness across all configurations, with further improvements when integrated with the ES modulator, confirming the effectiveness of the ES modulator in enhancing resilience against cross-device variations.
TABLE VII: Average Extraction Accuracy Under Varying Shooting Distances and Angles.
@@
96.28 97.69 97.24 95.27 94.61 97.67 95.33 95.00 94.67 93.67
Watermark Robustness (against Print-camera). Print-camera shooting is also a common method for capturing text content. We first print all the encoded text images on A4 paper at 600 dpi using the printer HP M132nw and take photos with a handheld cellphone HONOR 30S. The shooting distance is 20cm and the shooting angle is $0^{\circ}$. Subsequently, watermark extraction is performed on the images. As shown in Table V, CoreMark maintains stable high accuracy, clearly outperforming the best competing baseline, Yang23 [9]. When the Chinese font size exceeds 16pt, the $ACC$ of Yang23 surpasses 90%. But when the font size is smaller, its extraction accuracy drops sharply. To further assess the robustness of CoreMark against print-camera, we also evaluate it under other shooting conditions. The results, shown in Table VII, demonstrate consistently high accuracy across all conditions, confirming the strong robustness of CoreMark against common real-world perturbations such as distance variation and angular deviation during image capture.
@@
Watermark Generalizability (w.r.t. Different Fonts). We evaluate the robustness and imperceptibility of several commonly used fonts in both Chinese and English. The font sizes for Chinese and English fonts are set to 14pt and 18pt, respectively. For each font, we test 30 text images, and Table V records the detailed results. Under the Screenshot, CoreMark attains near-perfect accuracy, reaching up to 99.39% on English (TNR) and 98.19% on Chinese (DengXian). Under more challenging conditions such as Print-scan and Print-camera, CoreMark maintains robust performance, demonstrating its superior resilience to physical distortions. Yang23 performs better than Wu04, particularly on English fonts, but still falls short compared to CoreMark. These results validate the generalizability of CoreMark across diverse font styles.
Watermark Generalizability (w.r.t. Foreign Languages). We also evaluate the adaptability of CoreMark to languages beyond English and Chinese, including Arabic, French, Japanese, Korean, Russian, and Spanish. For each language, we test 30 text images. Table VI presents the PSNR, SSIM, and $ACC$ values of six languages. The results indicate that Arabic exhibits slightly lower robustness compared to other languages, but it achieves over 90% accuracy in print-scan and screenshot scenarios, and over 85% accuracy in resisting print-camera. Regardless of invisibility or robustness, CoreMark performs significantly better than the baseline methods.
Figure 13: Reducing examples in (a and b) Arabic, (c) and (d) Japanese, (e) and (f) Korean, (g) and (h) Russian.